# Project webpage: dino-tracker.github.io

### v1

@@ 3.3 Objective
where $\lambda_{*}$ sets the relative weights between the terms. We use a fixed set of $\lambda_{*}$ in all our experiments. See Appendix 0.A.2 for implementation and training details.
@@ 4 Results
[…]
@@ 4.2 Ablations and Analysis
[…]
@@ 5 Discussion and Conclusions
Regarding limitations, while our method excels in associating points across long-term occlusions, we do not model point trajectories behind occluders. Previous methods achieve this using synthetic data for supervision, or lifting tracking into 3D. Furthermore, our performance depends on the information encoded in DINO’s pre-trained features. We observed that in challenging videos for which there is almost no optical flow supervision and there are multiple semantically-similar objects, trajectories may jump from one object to another. This is because raw DINO is mostly dominated by semantic information.
@@ Acknowledgements
We would like to thank Rafail Fridman for his insightful remarks and assistance. We would also like to thank the authors of Omnimotion for providing the trained weights for TAP-Vid-DAVIS and TAP-Vid-Kinetics videos. The project was supported by an ERC starting grant OmniVideo (10111768), and the Carolito Stiftung.
@@ 0.A.6 Benchmarks Evaluation
Appendix 0.B Training time.

### v2

@@ 3.3 Objective
where $\lambda_{*}$ sets the relative weights between the terms. We use a fixed set of $\lambda_{*}$ in all our experiments. See Appendix 0.A.2 for further implementation details and Appendix 0.B for complexity details.
@@ 4 Results
Method DAVIS-256 DAVIS-480 Kinetics-256 Kinetics-480 BADJA $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{\textit{seg}}$ $\delta^{3px}$ RAFT [47] 56.7 66.7 50.4 60.5 45.0 5.8 DINOv2 [38] 61.4 64.7 60.3 61.0 62.8 8.4 TAP-Net [12] 53.4 81.4 38.4 66.4 79.0 46.0 61.7 86.6 48.5 67.1 81.5 47.7 45.4 9.6 PIPs++ [63] 71.5 73.6 68.2 70.8 59.0 9.8 TAPIR [13] 74.7 89.4 62.8 77.3 89.5 65.7 69.5 89.1 57.3 69.8 86.7 57.5 68.7 10.5 Co-Tracker [26] 79.2 89.3 65.1 79.4 89.5 65.6 72.9 88.9 59.9 72.8 88.9 59.8 64.0 11.2 Omnimotion [51] 67.5 85.3 51.7 74.1 84.5 58.4 69.2 89.2 55.0 45.2 6.9 Ours 78.2 87.5 62.3 80.4 88.1 64.6 73.3 88.5 59.7 74.3 89.2 60.9 72.4 14.3
– supervised.   – test-time training.
@@ 4.2 Ablations and Analysis
DINO features are the cornerstone of our framework. But which DINO features should we use? Tab. 3 shows track position accuracy for different choices of DINOv2 ViT-L/14 facets. Using tokens extracted from the 16th layer performs the best, and we use these DINO features in all experiments.
@@
Table 3: DINO’s feature layer ablation. We evaluate tracking performance using DINOv2 ViT-L/14 features extracted from different layers and facets. We report track position accuracy ($\delta^{x}_{avg}$) on TAP-Vid-DAVIS 480. Based on these results we use tokens extracted from the 16th layer.
@@ 5 Discussion and Conclusions
Regarding limitations, while our method excels in associating points across long-term occlusions, we do not model point trajectories behind occluders. Previous methods achieve this using synthetic data for supervision, or lifting tracking into 3D. However, a simple interpolation technique such as cubic spline can give plausible tracks during occlusion (see our SM website for examples). Furthermore, our performance depends on the information encoded in DINO’s pre-trained features. We observed that in challenging videos for which there is almost no optical flow supervision and there are multiple semantically-similar objects, trajectories may jump from one object to another. This is because raw DINO is mostly dominated by semantic information.
@@ Acknowledgements
We would like to thank Rafail Fridman for his insightful remarks and assistance. We would also like to thank the authors of Omnimotion for providing the trained weights for TAP-Vid-DAVIS and TAP-Vid-Kinetics videos. The project was supported by an ERC starting grant OmniVideo (10111768), by Shimon and Golde Picker, and by the Carolito Stiftung.
@@ 0.A.6 Benchmarks Evaluation
Appendix 0.B Complexity
0.B.1 Training time.
@@ Appendix 0.B Complexity
To improve training efficiency, we show that DINO-Tracker can be trained only on a subset of the frames, while still evaluating on all frames during inference. Tab. 4 reports our performance when training only $50\%$ or $25\%$ of the frames, thus reducing the training time by the same factor. As seen, our method maintains its performance when trained on $50\%$ and is competitive when trained on $25\%$ of the frames. This demonstrates the ability of our tracker to generalize to unseen frames and suggest it can be extended efficiently to longer videos.
Table 4: Generalization when training on every 2nd and every 4th frame in each video (TAP-Vid-DAVIS-480) on a single A100.
$\delta_{avg}^{x}$ AJ OA train time
Ours 80.7 65.3 88.5 90 min.
Ours, every 2nd frame 80.6 65.7 88.5 45 min.
Ours, every 4th frame 79.1 64.0 87.2 22.5 min.
0.B.2 Runtime and Memory
We measure the required compute for full inference (both position and visibility) of DINO-Tracker and feed-forward competitors. Tab. 5 reports average runtime and allocated memory on TAP-Vid-DAVIS-480 on a single A100 for our tracker and feed-forward methods. Most of our runtime is used for visibility prediction, yet once trained, our total inference time is fastest (note that PIPS++ cannot predict visibility) and is less memory-consuming than TAPIR.
Table 5: DAVIS-480 inference time and memory on a single A100.
Co-Tracker TAPIR PIPS++ Ours Ours
(full) (full) (pos. only) (pos. only) (full)
Time (sec) 345.8 110.4 8.6 4.3 80.5
GPU-Mem (GB) 19.2 60.0 12.0 15.2 52.6